My issue if finding a LCM to this equation. My teacher didn't explain it very well, however, I think AFTER obtaining the LCM I understand how to solve the equation. This particular problem is:
7/(x-3) - (2/7) = -10/(3x-9)

Answer by checkley77(12844)

7/(x-3) - (2/7) = -10/(3x-9)
[7*7-2(x-3)]/7(x-3)=-10/(3x-9)
[49-2x+6]/(7x-21)=-10/(3x-9)
[-2x+55]/(7x-21)=-10/(3x-9) cross multiply.
-10(7x-21)=(3x-9)(-2x+55)
-70x+210=-6x^2+18x+165x-495
-70x+210=-6x^2+183x-495
6x^2-70x-183x+210+495=0
6x^2-253x+705=0
(6X-235)(X-3)=0
6X-235=0
6X=235
X=235/6 ANS.
X-3=0
X=3 ANS.
PROOF:
7/(x-3) - (2/7) = -10/(3x-9)
7/[(235/6)-3]-2/7=-10/[(3*235/6)-9]
7/[(235-3*6)/6]-2/7=-10/[(705/6)-9]
7/[217/6]-2/7=-10/[705-9*6]/6
7*6/217-2/7=-10/[705-54]/6
42/217-2/7=-10/[651]/6
.1935-.2857=-10*6/651
-.0922=-60/651
-.0922=-.0922